OpenStudy (waheguru):

y = –x + 1 how to i convert this so i can graph it

OpenStudy (waheguru):

a+b+c=0 form

OpenStudy (hoblos):

x+y-1=0

OpenStudy (waheguru):

how do i graph it then

OpenStudy (anonymous):

Yeah a question! I would use the slope-intercept formula to graph. Begin by converting -1 into rise over run y = (-1/1)x + 1 Now to graph, begin with the y-intercept, placing a dot on (0,1) From (0,1) go down one and to the right one, that is (1,0) and connect these two dots.

OpenStudy (dumbcow):

it is already in "y=mx+b" form slope = -1 y-intercept = 1 go to (0,1) then go down 1 over 1 and make another point draw line through the points and there is your line

OpenStudy (hoblos):

you need two points to graph this line first replace x=0 => y=1 => first point is (0,1) then replace y=0 => x=1 => second point is (1,0) join these two points

OpenStudy (anonymous):

a + b + c = 0 is not a good form to graph from. At least not for me.

OpenStudy (waheguru):

thx